A 13 kg solid cylinder with a 54 cm diameter rolls without slipping down a 30 degree incline from a height of 1.25 meters. If a solid cylinder has a moment of inertia, I=½(MR2), what will its speed be if it rolls from a height of 1.25 meters down a 60-degree incline? this as a Free Respons Question!
Answer #1

using conservation of energy

pE at top = rotational kE + linear kE

m g h = 0.5* Iw^2 + 0.5 m v^2

mg h = 0.5* (0.5 m r^2) (v/r) ^2 + 0.5* m v^2

9.8* 1.25 = 0.25* v^2 + 0.5* v^2

v = 4.04 m/s

===========

final velocity at bottom is dependent only on height. Its independent of angle of inclination

=========

Comment in case any doubt, will reply for sure . Goodluck